# To Make, or to Take, That Is the Question: Impact of LOB Mechanics on Natural Trading Strategies

### v1

@@ 7.3 Results
Threshold Logistic Regression Random Forest # Fills / Day Avg Ret Stdev Holding Sharpe p-value # Fills / Day Avg Ret Stdev Holding Sharpe p-value 0.00 15380 $-0.44$ $2.73$ 11.39 $-109.00$ $0.5000$ 15380 $-0.44$ $2.73$ 11.39 $-109.00$ $0.5000$ 0.02 14266 $-0.45$ $2.81$ 12.28 $-99.98$ $0.6369$ 15066 $-0.43$ $2.68$ 12.54 $-106.78$ $0.3933$ 0.04 12650 $-0.46$ $2.97$ 13.85 $-92.82$ $0.8177$ 14626 $-0.43$ $2.71$ 12.92 $-105.31$ $0.4158$ 0.06 10972 $-0.45$ $3.20$ 15.97 $-74.99$ $0.6278$ 13868 $-0.44$ $2.77$ 13.62 $-99.97$ $0.5441$ 0.08 9364 $-0.44$ $3.46$ 18.71 $-67.29$ $0.5270$ 12856 $-0.45$ $2.87$ 14.69 $-98.61$ $0.6727$ 0.10 7780 $-0.44$ $3.77$ 22.52 $-56.42$ $0.5230$ 11700 $-0.44$ $2.97$ 16.14 $-96.94$ $0.5340$ 0.12 6332 $-0.44$ $4.24$ 27.67 $-45.39$ $0.5187$ 10354 $-0.44$ $3.17$ 18.24 $-91.83$ $0.5305$ 0.14 4956 $-0.43$ $4.72$ 35.35 $-37.41$ $0.4469$ 8918 $-0.45$ $3.42$ 21.18 $-75.09$ $0.5897$ 0.16 3598 $-0.43$ $5.74$ 48.68 $-29.20$ $0.4867$ 7630 $-0.41$ $3.68$ 24.75 $-63.28$ $0.2555$ 0.18 2494 $-0.40$ $6.72$ 70.24 $-23.20$ $0.3908$ 6374 $-0.41$ $4.07$ 29.63 $-55.68$ $0.3011$ 0.20 1644 $-0.39$ $8.10$ 106.53 $-21.26$ $0.3975$ 5224 $-0.36$ $4.54$ 36.15 $-42.96$ $0.1011$ 0.22 1060 $-0.22$ $10.63$ 165.32 $-15.88$ $0.2504$ 4146 $-0.39$ $5.17$ 45.54 $-42.23$ $0.2794$ 0.24 654 $+0.71$ $13.69$ 267.66 $+11.97$ $0.0160$ 3222 $-0.39$ $5.90$ 58.61 $-34.28$ $0.3252$ 0.26 426 $+1.21$ $18.10$ 409.26 $+26.12$ $0.0303$ 2460 $-0.32$ $6.80$ 76.73 $-24.85$ $0.1948$ 0.28 270 $+0.90$ $22.52$ 595.97 $+12.32$ $0.1650$ 1826 $-0.21$ $7.82$ 103.27 $-18.78$ $0.1076$ 0.30 - - - - - - 1330 $+0.03$ $8.90$ 141.75 $+1.30$ $0.0287$ 0.32 - - - - - - 972 $+0.31$ $9.65$ 194.22 $+12.08$ $0.0078$ 0.34 - - - - - - 692 $+0.59$ $12.20$ 272.14 $+13.48$ $0.0133$ 0.36 - - - - - - 454 $+0.02$ $14.45$ 414.03 $+0.30$ $0.2501$
Figure 17: PnL plot of the model-based MM strategy, net of fees.
Figure 18: Histogram of pre-fee returns of roundtrip trades of the model-based MM strategy.
@@ 7.4 Feature Importance
    Trade Volume Dynamics: In the recent past, the last trade returns should lack a strong autocorrelated trend, indicating that the price oscillates rather than trending strongly in one direction. Following such price action, there is a sudden drop in the top bid price (within the last 100 milliseconds) to a new local minimum, where there is minimal pre-existing liquidity.
  4. 4.
@@
  • The high importance score of the top feature, ret_autocov_sum_30s_w0, stresses the importance of recent price autocovariance as an explanatory variable. This corroborates the finding that negative autocovariance of recent returns (implying price oscillations up and down) are an important precursor of reversals, which we previously inferred from logistic regression’s large negative coefficients for ret_autocov_5s_w*, =0,1,2*=0,1,2∗ = 0 , 1 , 2.
@@
  • The logistic regression model applied a large negative coefficient to the feature sell_count_30s_w0, whose significance in predicting reversals is further corroborates by its high (normalized) permutation importance score of 0.27 in the random forest model.
@@
A concluding afterword notes that all of the explanatory variables with a high permutation score for the random forst model also exhibit a large (in absolute terms) coefficient in our logistic regression model; and conversely, there is no feature with a very large coefficient that does not emerge as one with a high permutation importance score. Thus, we conclude that, while the permutation importance analysis does not allow us to paint as detailed a picture as the coefficient analysis (since it only tells us how important a feature is without telling us how it influences the model’s predictions), its results are consistent with our earlier findings.
@@ 8 Conclusion
Through granular empirical analysis, we have described the effect of (i) basic order book mechanics and (ii) the strong persistence of price changes from the immediate to short timescales, revealing interactions between returns, queue sizes, and orders’ queue positions. We further examined the impact of these interactions on several natural (and commonly-cited) trading strategies, demonstrating that they are all highly unprofitable in practice. Finally, we used the understanding gained to identify situations (Reversals) where a successful maker strategy can operate by posting orders during an adverse imbalance and benefiting from an early queue position if market conditions shift favorably; we constructed a signal to detect these opportunities and demonstrated its efficacy.
In Section 1.1.2, we informally proposed an “Unprofitability Principle": forecasting ability and ease of exploitation are anti-correlated. This is really just an aspect of the set of ideas that make up the EMH, and it has been amply verified in our experiments. Even though we demonstrated an apparently profitable Reversal strategy, it required a relatively sophisticated analysis, offered relatively few trading opportunities, and its scalability is uncertain.
Many academic studies are bedeviled by one or both of the following difficulties. First, there is often a lack of access to data that provides the full picture: even with L3 data—the highest granularity generally available—it captures only what happened, not what did not happen (such as unfilled orders or cancellation attempts), which may be just as important. Second, studies frequently fail to examine the problem at a sufficiently granular level to uncover the influence of order book mechanics on outcomes. By conducting live trading experiments, we address both these challenges, leveraging a methodology that provides insights unattainable with other approaches.
Like all liquid LOB markets, the crypto exchange we used is highly competitive. The lack of apparent profitability for HFT is partly a consequence of the mechanics of an LOB market, partly a consequence of competition between traders, and pervasively a consequence of the exchange fees (the spread being very small). Indeed, given that the exchange takes 1 bp of every trade, it is hard to assess how Balance Inventory HFT, with its huge numbers of orders, could survive this extraction of wealth without a flow of Optimal Execution trading to fund it: it would be a negative-sum game.
In the end, the certain winners are, as so often, the service providers, in this case the exchanges. Perhaps the answer to the question in our title is “Neither a Maker nor a Taker be; instead, set up an Exchange."
@@ Appendix B Model Coefficients
Figure 30: Coefficients of the logistic regression model grouped by feature categories.
Momentum Feature Coefficient value ret_autocov_100ms_w0 -0.064670 ret_autocov_100ms_w1 -0.025752 ret_autocov_100ms_w2 -0.057328 ret_sum_100ms_w0 0.171598 ret_sum_100ms_w1 0.017349 ret_sum_100ms_w2 -0.070002 trade_intensity_100ms_w0 -0.028381 trade_intensity_100ms_w1 -0.048367 trade_intensity_100ms_w2 0.025111 ret_autocov_1s_w0 -0.156862 ret_autocov_1s_w1 0.002328 ret_autocov_1s_w2 -0.009059 ret_sum_1s_w0 -0.012417 ret_sum_1s_w1 0.006796 ret_sum_1s_w2 0.014250 trade_intensity_1s_w0 -0.027251 trade_intensity_1s_w1 -0.048310 trade_intensity_1s_w2 -0.039917 ret_autocov_5s_w0 -0.433133 ret_autocov_5s_w1 -0.261970 ret_autocov_5s_w2 -0.317830 ret_sum_5s_w0 -0.070847 ret_sum_5s_w1 0.016905 ret_sum_5s_w2 -0.076009 trade_intensity_5s_w0 -0.066931 trade_intensity_5s_w1 -0.044181 trade_intensity_5s_w2 -0.056927 ret_autocov_30s_w0 0.111506 ret_autocov_30s_w1 -0.044950 ret_autocov_30s_w2 -0.102125 ret_sum_30s_w0 -0.020848 ret_sum_30s_w1 0.033951 ret_sum_30s_w2 -0.038579 trade_intensity_30s_w0 -0.310782 trade_intensity_30s_w1 -0.318268 trade_intensity_30s_w2 -0.100313 ret_autocov_300s_w0 0.055847 ret_autocov_300s_w1 -0.105477 ret_autocov_300s_w2 -0.015074 ret_sum_300s_w0 -0.085791 ret_sum_300s_w1 0.048547 ret_sum_300s_w2 0.010486 trade_intensity_300s_w0 -0.013885 trade_intensity_300s_w1 -0.012627 trade_intensity_300s_w2 -0.012789 Order Book Feature Coefficient value age -0.056527 ob_ask_liq 0.001471 ob_bid_liq -0.187389 ob_half -0.049546 ob_other_half -0.007837 totb_mean -0.198196 Price Dynamics Feature Coefficient value amplitude_100ms_w0 -0.134841 amplitude_100ms_w1 -0.002490 amplitude_100ms_w2 -0.015208 ret_vwap_100ms_w0 0.054331 ret_vwap_100ms_w1 -0.106379 ret_vwap_100ms_w2 0.035287 amplitude_1s_w0 0.009782 amplitude_1s_w1 0.007194 amplitude_1s_w2 0.019774 ret_vwap_1s_w0 -0.007336 ret_vwap_1s_w1 -0.019845 ret_vwap_1s_w2 0.020953 amplitude_5s_w0 0.103106 amplitude_5s_w1 0.051163 amplitude_5s_w2 0.101474 ret_vwap_5s_w0 0.041889 ret_vwap_5s_w1 0.010773 ret_vwap_5s_w2 -0.009981 amplitude_30s_w0 -0.025883 amplitude_30s_w1 0.009798 amplitude_30s_w2 0.049859 ret_vwap_30s_w0 -0.051273 ret_vwap_30s_w1 -0.058798 ret_vwap_30s_w2 0.019591 amplitude_300s_w0 0.041433 amplitude_300s_w1 0.023701 amplitude_300s_w2 -0.009184 ret_vwap_300s_w0 0.033777 ret_vwap_300s_w1 -0.010073 ret_vwap_300s_w2 0.017873 stdev_100 -0.014769 stdev_500 0.069001
Trade Volume Patterns Feature Coefficient value max_size_100ms_w0 -0.015398 max_size_100ms_w1 -0.064591 max_size_100ms_w2 0.001329 avg_size_100ms_w0 -0.012507 avg_size_100ms_w1 0.031008 avg_size_100ms_w2 0.004749 buy_count_100ms_w0 -0.002806 buy_count_100ms_w1 -0.028945 buy_count_100ms_w2 0.044143 sell_count_100ms_w0 -0.042881 sell_count_100ms_w1 -0.031871 sell_count_100ms_w2 -0.010898 total_buy_100ms_w0 0.017631 total_buy_100ms_w1 0.031328 total_buy_100ms_w2 -0.025174 total_sell_100ms_w0 0.027392 total_sell_100ms_w1 0.037905 total_sell_100ms_w2 0.019188 max_size_1s_w0 0.009413 max_size_1s_w1 -0.024575 max_size_1s_w2 -0.006610 avg_size_1s_w0 0.021593 avg_size_1s_w1 0.015751 avg_size_1s_w2 0.006351 buy_count_1s_w0 -0.034406 buy_count_1s_w1 -0.046010 buy_count_1s_w2 -0.023821 sell_count_1s_w0 0.094811 sell_count_1s_w1 0.024451 sell_count_1s_w2 -0.000842 total_buy_1s_w0 0.002977 total_buy_1s_w1 -0.003422 total_buy_1s_w2 0.004587 total_sell_1s_w0 -0.049719 total_sell_1s_w1 -0.012063 total_sell_1s_w2 -0.011760 max_size_5s_w0 0.005391 max_size_5s_w1 0.023364 max_size_5s_w2 0.000746 avg_size_5s_w0 0.000672 avg_size_5s_w1 -0.002132 avg_size_5s_w2 0.019611 buy_count_5s_w0 0.091383 buy_count_5s_w1 -0.026798 buy_count_5s_w2 -0.019723 sell_count_5s_w0 0.077718 sell_count_5s_w1 0.094188 sell_count_5s_w2 0.103399 total_buy_5s_w0 -0.085354 total_buy_5s_w1 -0.041604 total_buy_5s_w2 0.016422 total_sell_5s_w0 0.023385 total_sell_5s_w1 -0.077652 total_sell_5s_w2 -0.068970 max_size_30s_w0 -0.053188 max_size_30s_w1 -0.004395 max_size_30s_w2 -0.009718 avg_size_30s_w0 0.002545 avg_size_30s_w1 -0.010833 avg_size_30s_w2 -0.033334 buy_count_30s_w0 0.021358 buy_count_30s_w1 -0.046941 buy_count_30s_w2 -0.019813 sell_count_30s_w0 -0.172218 sell_count_30s_w1 -0.098570 sell_count_30s_w2 0.081075 total_buy_30s_w0 0.023254 total_buy_30s_w1 0.070232 total_buy_30s_w2 0.021899 total_sell_30s_w0 0.170826 total_sell_30s_w1 0.001023 total_sell_30s_w2 -0.080049 max_size_300s_w0 -0.004522 max_size_300s_w1 -0.024765 max_size_300s_w2 -0.026390 avg_size_300s_w0 0.092860 avg_size_300s_w1 -0.006021 avg_size_300s_w2 -0.019310 buy_count_300s_w0 -0.077479 buy_count_300s_w1 0.124425 buy_count_300s_w2 0.017996 sell_count_300s_w0 0.021356 sell_count_300s_w1 -0.036636 sell_count_300s_w2 -0.148054 total_buy_300s_w0 0.081260 total_buy_300s_w1 -0.129170 total_buy_300s_w2 -0.002067 total_sell_300s_w0 0.047449 total_sell_300s_w1 -0.008577 total_sell_300s_w2 0.147168

### v2

@@ 7.3  Results
Threshold Logistic Regression Random Forest # Fills / Day Avg Ret Stdev Holding Sharpe p-value # Fills / Day Avg Ret Stdev Holding Sharpe p-value 0.00 15380 0.44-0.44 $2.73$ 11.39 109.00-109.00 $0.5000$ 15380 0.44-0.44 $2.73$ 11.39 109.00-109.00 $0.5000$ 0.02 14266 0.45-0.45 $2.81$ 12.28 99.98-99.98 $0.6369$ 15066 0.43-0.43 $2.68$ 12.54 106.78-106.78 $0.3933$ 0.04 12650 0.46-0.46 $2.97$ 13.85 92.82-92.82 $0.8177$ 14626 0.43-0.43 $2.71$ 12.92 105.31-105.31 $0.4158$ 0.06 10972 0.45-0.45 $3.20$ 15.97 74.99-74.99 $0.6278$ 13868 0.44-0.44 $2.77$ 13.62 99.97-99.97 $0.5441$ 0.08 9364 0.44-0.44 $3.46$ 18.71 67.29-67.29 $0.5270$ 12856 0.45-0.45 $2.87$ 14.69 98.61-98.61 $0.6727$ 0.10 7780 0.44-0.44 $3.77$ 22.52 56.42-56.42 $0.5230$ 11700 0.44-0.44 $2.97$ 16.14 96.94-96.94 $0.5340$ 0.12 6332 0.44-0.44 $4.24$ 27.67 45.39-45.39 $0.5187$ 10354 0.44-0.44 $3.17$ 18.24 91.83-91.83 $0.5305$ 0.14 4956 0.43-0.43 $4.72$ 35.35 37.41-37.41 $0.4469$ 8918 0.45-0.45 $3.42$ 21.18 75.09-75.09 $0.5897$ 0.16 3598 0.43-0.43 $5.74$ 48.68 29.20-29.20 $0.4867$ 7630 0.41-0.41 $3.68$ 24.75 63.28-63.28 $0.2555$ 0.18 2494 0.40-0.40 $6.72$ 70.24 23.20-23.20 $0.3908$ 6374 0.41-0.41 $4.07$ 29.63 55.68-55.68 $0.3011$ 0.20 1644 0.39-0.39 $8.10$ 106.53 21.26-21.26 $0.3975$ 5224 0.36-0.36 $4.54$ 36.15 42.96-42.96 $0.1011$ 0.22 1060 0.22-0.22 $10.63$ 165.32 15.88-15.88 $0.2504$ 4146 0.39-0.39 $5.17$ 45.54 42.23-42.23 $0.2794$ 0.24 654 +0.71+0.71 $13.69$ 267.66 +11.97+11.97 $0.0160$ 3222 0.39-0.39 $5.90$ 58.61 34.28-34.28 $0.3252$ 0.26 426 +1.21+1.21 $18.10$ 409.26 +26.12+26.12 $0.0303$ 2460 0.32-0.32 $6.80$ 76.73 24.85-24.85 $0.1948$ 0.28 270 +0.90+0.90 $22.52$ 595.97 +12.32+12.32 $0.1650$ 1826 0.21-0.21 $7.82$ 103.27 18.78-18.78 $0.1076$ 0.30 - - - - - - 1330 +0.03+0.03 $8.90$ 141.75 +1.30+1.30 $0.0287$ 0.32 - - - - - - 972 +0.31+0.31 $9.65$ 194.22 +12.08+12.08 $0.0078$ 0.34 - - - - - - 692 +0.59+0.59 $12.20$ 272.14 +13.48+13.48 $0.0133$ 0.36 - - - - - - 454 +0.02+0.02 $14.45$ 414.03 +0.30+0.30 $0.2501$
Figure 17: PnL plot of the model-based $M$ strategy, net of fees.
Figure 18: Histogram of pre-fee returns of roundtrip trades of the model-based $M$ strategy.
@@ 7.4  Feature Importance
  • The high importance score of the top feature, ret_autocov_sum_30s_w0, stresses the importance of recent price autocovariance as an explanatory variable. This corroborates the finding that negative autocovariance of recent returns (implying price oscillations up and down) are an important precursor of reversals, which we previously inferred from logistic regression’s large negative coefficients for ret_autocov_5s_w*, =0,1,2*=0,1,2.
@@
  • The logistic regression model applied a large negative coefficient to the feature sell_count_30s_w0, whose significance in predicting reversals is further corroborated by its high (normalized) permutation importance score of 0.27 in the random forest model.
@@
A concluding afterword notes that all of the explanatory variables with a high permutation score for the random forest model also exhibit a large (in absolute terms) coefficient in our logistic regression model; and conversely, there is no feature with a very large coefficient that does not emerge as one with a high permutation importance score. Thus, we conclude that, while the permutation importance analysis does not allow us to paint as detailed a picture as the coefficient analysis (since it only tells us how important a feature is without telling us how it influences the model’s predictions), its results are consistent with our earlier findings.
@@ 8  Conclusion
Through granular empirical analysis, we have described the effect of (i) basic order book mechanics and (ii) the strong persistence of price changes from the immediate to short timescales, revealing interactions between returns, queue sizes, and orders’ queue positions. These interactions lead to negative correlation between an order’s fill probability and its post-fill returns, creating a fundamental challenge in profitable maker trading at the touch. The practical consequences in high-frequency trading are immense, rendering several natural (and commonly-cited) trading strategies highly unprofitable in practice.
To overcome these challenges, a key to successful market making is to avoid being adversely selected. While this can in principle be achieved ”defensively” by cancelling orders when adverse selection becomes likely (itself a complex task), our work explores a more proactive strategy: developing counter-signals to the order book imbalance. Such signals allow for the submission of orders during an apparently adverse imbalance to benefit from an early queue position when conditions shift favorably. To show that this can be done, we develop such a signal and demonstrate its efficacy: it significantly outperforms benchmarks and gives an indication of the type of features that might be helpful in predicting when the order book imbalance falsely predicts the next price change; the strategy is slightly profitable with min-sized orders and assuming low fees, albeit those profits are likely not scalable; however, we hope that our approach can give practitioners a pointer on what direction to focus their efforts. The specific model we present is one of many possible constructions, intended to demonstrate a proof-of-concept rather than to serve as an optimized, production-ready strategy.
Our model’s performance, which required a complex feature set to achieve even modest (and likely non-scalable) profits, affirms the ”Unprofitability Principle” we informally introduced earlier as: forecasting ability and ease of exploitation are anti-correlated. This principle helps explain why so many academic studies struggle to produce practically viable results—a tendency exemplified by a research focus on fill probabilities alone, which, by ignoring their negative correlation with returns, risks optimizing for a pyrrhic victory. This flawed approach is symptomatic of two broader challenges that often bedevil academic research: (i) a failure to examine the problem at a sufficiently granular level to uncover the influence of order book mechanics on outcomes, and (ii) a lack of access to data that provides the full picture, as even L3 data show only what happened, not what did not happen (such as failed cancellations), which can be equally important. By conducting live trading experiments, we address both these challenges, leveraging a methodology that provides insights difficult to attain with other approaches.
Ultimately, like all liquid LOB markets, the crypto exchange we used is highly competitive. The pervasive barrier to profitability for HFT (maker or taker) is partly a consequence of the mechanics of an LOB market, partly a consequence of competition between traders, and pervasively a consequence of the persistent wealth extraction via exchange fees.
Funding
AYS acknowledges support from NSERC Discovery Grant RGPIN-2024-05996.
Data Availability
The data that support the findings of this study were collected by the authors and are not publicly available.
Declaration of Competing Interests
The authors declare no competing financial or personal interests in relation to the work described.
@@ Appendix B Model Coefficients
Table 5: Coefficients of the logistic regression model grouped by feature categories.
Feature Coefficient value
Momentum
ret_autocov_100ms_w0 -0.064670
ret_autocov_100ms_w1 -0.025752
ret_autocov_100ms_w2 -0.057328
ret_sum_100ms_w0 0.171598
ret_sum_100ms_w1 0.017349
ret_sum_100ms_w2 -0.070002
trade_intensity_100ms_w0 -0.028381
trade_intensity_100ms_w1 -0.048367
trade_intensity_100ms_w2 0.025111
ret_autocov_1s_w0 -0.156862
ret_autocov_1s_w1 0.002328
ret_autocov_1s_w2 -0.009059
ret_sum_1s_w0 -0.012417
ret_sum_1s_w1 0.006796
ret_sum_1s_w2 0.014250
trade_intensity_1s_w0 -0.027251
trade_intensity_1s_w1 -0.048310
trade_intensity_1s_w2 -0.039917
ret_autocov_5s_w0 -0.433133
ret_autocov_5s_w1 -0.261970
ret_autocov_5s_w2 -0.317830
ret_sum_5s_w0 -0.070847
ret_sum_5s_w1 0.016905
ret_sum_5s_w2 -0.076009
trade_intensity_5s_w0 -0.066931
trade_intensity_5s_w1 -0.044181
trade_intensity_5s_w2 -0.056927
ret_autocov_30s_w0 0.111506
ret_autocov_30s_w1 -0.044950
ret_autocov_30s_w2 -0.102125
ret_sum_30s_w0 -0.020848
ret_sum_30s_w1 0.033951
ret_sum_30s_w2 -0.038579
trade_intensity_30s_w0 -0.310782
trade_intensity_30s_w1 -0.318268
trade_intensity_30s_w2 -0.100313
ret_autocov_300s_w0 0.055847
ret_autocov_300s_w1 -0.105477
ret_autocov_300s_w2 -0.015074
ret_sum_300s_w0 -0.085791
ret_sum_300s_w1 0.048547
ret_sum_300s_w2 0.010486
trade_intensity_300s_w0 -0.013885
trade_intensity_300s_w1 -0.012627
trade_intensity_300s_w2 -0.012789
Order Book
age -0.056527
ob_ask_liq 0.001471
ob_bid_liq -0.187389
ob_half -0.049546
ob_other_half -0.007837
totb_mean -0.198196
Price Dynamics
amplitude_100ms_w0 -0.134841
amplitude_100ms_w1 -0.002490
amplitude_100ms_w2 -0.015208
ret_vwap_100ms_w0 0.054331
ret_vwap_100ms_w1 -0.106379
ret_vwap_100ms_w2 0.035287
amplitude_1s_w0 0.009782
amplitude_1s_w1 0.007194
amplitude_1s_w2 0.019774
ret_vwap_1s_w0 -0.007336
ret_vwap_1s_w1 -0.019845
ret_vwap_1s_w2 0.020953
amplitude_5s_w0 0.103106
amplitude_5s_w1 0.051163
amplitude_5s_w2 0.101474
ret_vwap_5s_w0 0.041889
ret_vwap_5s_w1 0.010773
ret_vwap_5s_w2 -0.009981
amplitude_30s_w0 -0.025883
amplitude_30s_w1 0.009798
amplitude_30s_w2 0.049859
ret_vwap_30s_w0 -0.051273
ret_vwap_30s_w1 -0.058798
ret_vwap_30s_w2 0.019591
amplitude_300s_w0 0.041433
amplitude_300s_w1 0.023701
amplitude_300s_w2 -0.009184
ret_vwap_300s_w0 0.033777
ret_vwap_300s_w1 -0.010073
ret_vwap_300s_w2 0.017873
stdev_100 -0.014769
stdev_500 0.069001
Trade Volume Patterns
max_size_100ms_w0 -0.015398
max_size_100ms_w1 -0.064591
max_size_100ms_w2 0.001329
avg_size_100ms_w0 -0.012507
avg_size_100ms_w1 0.031008
avg_size_100ms_w2 0.004749
buy_count_100ms_w0 -0.002806
buy_count_100ms_w1 -0.028945
buy_count_100ms_w2 0.044143
sell_count_100ms_w0 -0.042881
sell_count_100ms_w1 -0.031871
sell_count_100ms_w2 -0.010898
total_buy_100ms_w0 0.017631
total_buy_100ms_w1 0.031328
total_buy_100ms_w2 -0.025174
total_sell_100ms_w0 0.027392
total_sell_100ms_w1 0.037905
total_sell_100ms_w2 0.019188
max_size_1s_w0 0.009413
max_size_1s_w1 -0.024575
max_size_1s_w2 -0.006610
avg_size_1s_w0 0.021593
avg_size_1s_w1 0.015751
avg_size_1s_w2 0.006351
buy_count_1s_w0 -0.034406
buy_count_1s_w1 -0.046010
buy_count_1s_w2 -0.023821
sell_count_1s_w0 0.094811
sell_count_1s_w1 0.024451
sell_count_1s_w2 -0.000842
total_buy_1s_w0 0.002977
total_buy_1s_w1 -0.003422
total_buy_1s_w2 0.004587
total_sell_1s_w0 -0.049719
total_sell_1s_w1 -0.012063
total_sell_1s_w2 -0.011760
max_size_5s_w0 0.005391
max_size_5s_w1 0.023364
max_size_5s_w2 0.000746
avg_size_5s_w0 0.000672
avg_size_5s_w1 -0.002132
avg_size_5s_w2 0.019611
buy_count_5s_w0 0.091383
buy_count_5s_w1 -0.026798
buy_count_5s_w2 -0.019723
sell_count_5s_w0 0.077718
sell_count_5s_w1 0.094188
sell_count_5s_w2 0.103399
total_buy_5s_w0 -0.085354
total_buy_5s_w1 -0.041604
total_buy_5s_w2 0.016422
total_sell_5s_w0 0.023385
total_sell_5s_w1 -0.077652
total_sell_5s_w2 -0.068970
max_size_30s_w0 -0.053188
max_size_30s_w1 -0.004395
max_size_30s_w2 -0.009718
avg_size_30s_w0 0.002545
avg_size_30s_w1 -0.010833
avg_size_30s_w2 -0.033334
buy_count_30s_w0 0.021358
buy_count_30s_w1 -0.046941
buy_count_30s_w2 -0.019813
sell_count_30s_w0 -0.172218
sell_count_30s_w1 -0.098570
sell_count_30s_w2 0.081075
total_buy_30s_w0 0.023254
total_buy_30s_w1 0.070232
total_buy_30s_w2 0.021899
total_sell_30s_w0 0.170826
total_sell_30s_w1 0.001023
total_sell_30s_w2 -0.080049
max_size_300s_w0 -0.004522
max_size_300s_w1 -0.024765
max_size_300s_w2 -0.026390
avg_size_300s_w0 0.092860
avg_size_300s_w1 -0.006021
avg_size_300s_w2 -0.019310
buy_count_300s_w0 -0.077479
buy_count_300s_w1 0.124425
buy_count_300s_w2 0.017996
sell_count_300s_w0 0.021356
sell_count_300s_w1 -0.036636
sell_count_300s_w2 -0.148054
total_buy_300s_w0 0.081260
total_buy_300s_w1 -0.129170
total_buy_300s_w2 -0.002067
total_sell_300s_w0 0.047449
total_sell_300s_w1 -0.008577
total_sell_300s_w2 0.147168
Table 5: *